The product of two positive consecutive integers is 41 more their sum. Find the intergers.
Answer by unlockmath(1688)

Hello,
The key to these is understanding the words. Let's represent x and x+1 as the two positive consecutive integers.
Now we can set up an equation.
X(X+1)=X+(X+1)+41 rewrite it as:
x^2+x=2x+42 Subtract 2x and 42 from both sides and we get:
x^2-x-42=0 Now we can factor it to:
(x-7)(x+6)=0 This gives us:
x=7
x=-6 Since the numbers are positive x=7 is the answer. Of course the other integer is 8. Plug in 7 to the original equation and it will work out.
RJ
www.math-unlock.com